OpenStudy (ellecullen):

sin(a - b) = sina cosb + cosa sinb for all values of a and b. True or false?

OpenStudy (anonymous):

false, because the rules for sine a+b identities are. \[\sin(a +b) = \sin(a) ~\cos(b) + \cos(a) ~\sin(b)\]\[\sin(a -b) = \sin(a) ~\cos(b) - \cos(a) ~\sin(b)\]

OpenStudy (anonymous):

you have two of them mixed, and two sides are for sure not always equal.

OpenStudy (ellecullen):

Thank you :)

OpenStudy (anonymous):

\(\normalsize\color{black}{ \bf You}\) \(\normalsize\color{black}{ \bf welcome~!}\)